9. Factor completely. 5a2 – 125
A) 5(a – 5)2 B) 5a(a – 25) C) (5a – 1)(a – 125) D) 5(a + 5)(a – 5)

Answer by venugopalramana(3286)

9. Factor completely. 5a2 – 125
=5(A^2-25)=5(A^2-5^2)=5(A+5)(A-5).............D IS OK
A) 5(a – 5)2 B) 5a(a – 25) C) (5a – 1)(a – 125) D) 5(a + 5)(a – 5)